40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14633     Entry time: Thu May 23 10:18:39 2019
Author: Kruthi 
Type: Update 
Category: Cameras 
Subject: CCD calibration 

On Tuesday, I tried reproducing Pooja's measurements (https://nodus.ligo.caltech.edu:8081/40m/13986). The table below shows the values I got. Pictures of LED circuit, schematic and the setup are attached. The powermeter readings fluctuated quite a bit for input volatges (Vcc) > 8V, therefore, I expect a maximum uncertainity of 50µW to be on a safer side. Though the readings at lower input voltages didn't vary much over time (variation < 2µW), I don't know how relaible the Ophir powermeter is at such low power levels. The optical power output of LED was linear for input voltages 10V to 20V. I'll proceed with the CCD calibration soon.

Input Voltage (Vcc) in volts Optical power
0 (dark reading) 1.6 nW
2 55.4 µW
4 215.9 µW
6 0.398 mW
8 0.585 mW
10 0.769 mW
12 0.929 mW
14 1.065 mW
16 1.216 mW
18 1.330 mW
20 1.437 mW
22 1.484 mW
24 1.565 mW
26 1.644 mW
28 1.678 mW

Attachment 1: setup.jpeg  1.210 MB  Uploaded Thu May 23 12:29:10 2019  | Hide | Hide all
setup.jpeg
Attachment 2: led_circuit.jpeg  4.524 MB  Uploaded Thu May 23 12:39:11 2019  | Hide | Hide all
led_circuit.jpeg
Attachment 3: led_schematic.pdf  9 kB  Uploaded Thu May 23 12:39:35 2019  | Hide | Hide all
led_schematic.pdf
ELOG V3.1.3-